OpenStudy (anonymous):

In Mrs. Keller's class, the number of boys is five more than the number of girls. Let y represent the number of boys. Let x represent the number of girls. Which equation shows the relationship between y and x? y=x-5 y=5x y+x=5 y=x+5

OpenStudy (aripotta):

which one are you leaning towards?

OpenStudy (anonymous):

y=x-5

OpenStudy (aripotta):

y = x - 5 means the number of boys equals the number of girls, minus 5. the beginning of the question said it would be the number of girls PLUS 5. so the answer is y = x + 5